# Variational Tail Bounds for Norms of Random Vectors and Matrices

### v3

@@ 3 Simplified Bounds via Pushforward of Gaussians
We can recover the bound (9) in Example˜1 from Theorem˜1 up to a factor of order $\sqrt{\log c_{k}}$ that can be ignored if $c_{k}$ is not large relative to $\pi_{k}^{-1}$. The parameters of the problem in this example are $\eta_{1}=\sqrt{2}$, $\eta_{2}=0$, and $\Sigma=I$. By taking $\kappa_{0}\to\infty$, we have $f_{0}(x)=\rho_{0}{\left\lVert x\right\rVert}$ which corresponds to the choice
we have
$\geq 2\pi_{k}\,$,
using which we obtain the bound
$\overline{\nu}_{\Sigma}(\infty,2c_{k})$ $\leq\frac{c_{k}}{\pi_{k}}\,.$
Therefore, the bound (12) provided by Theorem˜1 implies
${\left\lVert X\right\rVert}$ $\leq 2\max_{i\in[N]}{\left\lVert u_{i}\right\rVert}^{2}_{2}\inf_{\begin{subarray}{c}p\geq 1\\ k\in[N]\end{subarray}}e^{t/(2p)}\sqrt{2p}c_{k}\left(\frac{c_{k}}{\pi_{k}}\right)^{1/(2p)}$
Let $k_{\star}\in[N]$ be the minimizer of $\min_{k\in[N]}c_{k}\max\left\{1,\sqrt{t-\log\pi_{k}}\right\}$ in (9). If we have $\log c_{k_{\star}}\lesssim t-\log\pi_{k_{\star}}$, our derivations confirm that, up to constant factors, Theorem˜1 recovers the bound (9) of Example˜1.
Next we will show how Example˜2 can be recovered from Theorem˜1. Again we have $\eta_{1}=\sqrt{2}$ and $\eta_{2}=0$. Also, recall that in this example $\mathcal{B}$ and $\mathcal{B}_{*}$ coincide with the unit $\ell_{2}$-ball. By taking $\rho_{0}\to+\infty$ we have $f_{0}(x)=\kappa_{0}{\left\lVert x\right\rVert}_{\Sigma^{-1/2}}^{2}/2$, and thus
$U=\nabla f_{0}(\Sigma^{1/2}G)$ $=\kappa_{0}G\,.$
$\Omega_{\Sigma}(p,\kappa_{0},\infty)$ $=\sqrt{2p}\kappa_{0}\left(\sqrt{\operatorname{tr}(\Sigma)}+\sqrt{2p{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}}\right)\,.$
$=\frac{\kappa_{0}}{\sqrt{2\pi}}e^{-\kappa_{0}^{-2}}\,$
Therefore, the bound (12) of Theorem˜1 implies
${\left\lVert X\right\rVert}={\left\lVert X\right\rVert}_{2}$ $\leq\inf_{\begin{subarray}{c}p\geq 1\\ \kappa_{0}>0\end{subarray}}e^{t/(2p)}\sqrt{2p}\kappa_{0}\left(\sqrt{\operatorname{tr}(\Sigma)}+\sqrt{2p{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}}\right)8^{1/(2p)}e^{1/(2p\kappa_{0}^{2})}$
$=\sqrt{2e}\inf_{p\geq 1}e^{(t+\log 8)/(2p)}\left(\sqrt{\operatorname{tr}(\Sigma)}+\sqrt{2p{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}}\right)$
$\leq 4\sqrt{e}\left(\sqrt{3\operatorname{tr}(\Sigma)}+\sqrt{\frac{1}{\log 2}{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}t}\right)\,$,
where the third line follows by evaluating the argument of the infimum at $p=1+t/\log 8$ and the fact that ${\left\lVert\Sigma\right\rVert}_{\mathrm{op}}\leq\operatorname{tr}(\Sigma)$. Therefore, up to constant factors, we have recovered the same bound on ${\left\lVert X\right\rVert}_{2}$ in Example˜2 through Theorem˜1.
4 A Concentration Inequality for Sum of Random Positive Semidefinite Matrices
We mentioned previously that we may easily adapt Lemma˜1 to apply under different moment conditions. Furthermore, the special structure of the norm of interest $\left\lVert\cdot\right\rVert$} can be leveraged to adjust the choice of $\operatorname{\mathbb{P}}_{0}$ in Lemma˜1. As an example, we use Lemma˜1 to recover, up to the constant factors, a result of [Zhi24, Theorem 1 ] that itself generalizes a concentration inequality due to [KL17, Theorem 9 ] for sample covariance matrices of sub-Gaussian random vectors. [^5] It has to be mentioned that the emphasis in [KL17] is on such concentration inequalities in possibly infinite dimensional spaces.
@@ 4 A Concentration Inequality for Sum of Random Positive Semidefinite Matrices
Since $\sum_{i\in[n]}\varepsilon_{i}u^{\scriptscriptstyle\textup{{T}}}Z_{i}u$ is a sum of independent symmetric random variables, it follows from a moment bound due to [Lat97, Corollary 2 ] that
@@
Next, we bound $\operatorname{\mathbb{P}}\left(\kappa_{0}^{2}\left|G^{\scriptscriptstyle\textup{{T}}}MG\right|\geq 1\right)$ for $M\in\mathbb{M}_{\mathrm{sym}}^{\ell}$ with ${\left\lVert M\right\rVert}_{\mathrm{op}}=1$. We use a simple form of a general technique due to [Cha19, Lemma 1.2 ]. Let $\widetilde{G}=(I-vv^{\scriptscriptstyle\textup{{T}}})G+\widetilde{g}v$ where $v$ is the singular vector of $M$ corresponding to the singular value 11, and $\widetilde{g}\sim\mathrm{Normal}(0,1)$ is independent of $G$. Since $G$ and $\widetilde{G}$ are both standard Gaussian random vectors, we have
@@
5 A Concentration Inequality for the Sample Covariance Matrix of Sub-exponential Random Vectors
[Zhi24, Theorem 3 ] also provides a concentration inequality for the sample covariance matrix of centered log-concave distributions. Following [ALPT09], [Zhi24] approximates the quadratic form induced by the sample covariance matrix by a term obtained through a certain direction-dependent truncation, and a term formed by the “peaky” residuals. The PAC-Bayesian argument is then used to obtain the tail bound for the “truncated” term, whereas the tail bound for peaky term is obtained using the decoupling-chaining method [Tal14, Proposition 9.4.2] as cited in [Zhi24], combined with the tail bound for the Euclidean norm of log-concave random vectors due to [Pao06, Thoerem 1.1 ] (see also [ALL+14, Theorem 2]). In fact, Paouris’s inequality is crucial in the proof of [Zhi24, Theorem 3 ] in the sense that it is tailored to log-concave measures instead of merely using the sub-exponential property. [^6] Log-concave measures are “regular” (see, e.g., [LW08, Definition 3.4], [Lat17, the statement following Definition 2], [Mur22]) which also implies that they are sub-exponential. Our next result, Theorem˜3, provides a concentration inequality for the sample covariance matrix of random vectors with sub-exponential distribution. While not adapted for log-concave distributions, without resorting to truncation techniques and solely relying on Lemma˜1, we can partially recover the result of [Zhi24, Theorem 3 ] with a matching confidence level. Specifically, under the extra condition that the number of samples is at least the cube of the effective rank of the population covariance matrix, the tail bound we provide below in Theorem˜3 is equivalent to the tail bound provided by [Zhi24, Theorem 3 ].
@@ 5 A Concentration Inequality for the Sample Covariance Matrix of Sub-exponential Random Vectors
The proof is similar to the proof of Theorem˜2. Let $X=\sum_{i=1}^{n}\left(Z_{i}Z_{i}^{\scriptscriptstyle\textup{{T}}}-\Sigma\right)$ and $U=\kappa_{0}^{2}GG^{\scriptscriptstyle\textup{{T}}}$ for a parameter $\kappa_{0}>0$ and a standard Gaussian random vector $G\in\mathbb{R}^{d}$ that is independent of everything else. We can apply Lemma˜1 as in the proof of Theorem˜2; modification is only required to approximate the moments of $\left|{\langle U,X\rangle}\right|$. Using the standard symmetrization argument, for any $u\in\mathbb{R}^{d}$ we have
@@
where $\varepsilon_{1},\dotsc,\varepsilon_{n}$ are Rademacher random variables independent of each other and everything else. It then follows from [Lat97, Corollary 2] that for $p\geq 2$ we have
@@
We can again show (20) still holds following the same argument used in the proof of Theorem˜2. Therefore, we have

### v4

@@ 3 Simplified Bounds via Pushforward of Gaussians
4 Applications to Random Matrices
4.1 A concentration inequality for sum of random positive semidefinite matrices
We mentioned previously that we may easily adapt Lemma˜1 to apply under different moment conditions. Furthermore, the special structure of the norm of interest $\left\lVert\cdot\right\rVert$} can be leveraged to adjust the choice of $\operatorname{\mathbb{P}}_{0}$ in Lemma˜1. As an example, we use Lemma˜1 to recover, up to the constant factors, a result of [56, Theorem 1 ] that itself generalizes a concentration inequality due to [23, Theorem 9 ] for sample covariance matrices of sub-Gaussian random vectors. [^4] The emphasis in [23] is on the concentration inequalities in possibly infinite dimensional spaces.
@@ 4.1 A concentration inequality for sum of random positive semidefinite matrices
Since $\sum_{i\in[n]}\varepsilon_{i}u^{\scriptscriptstyle\textup{{T}}}Z_{i}u$ is a sum of independent symmetric random variables, it follows from a moment bound due to [27, Corollary 2 ] that
@@
Next, we bound $\operatorname{\mathbb{P}}\left(\kappa_{0}^{2}\left|G^{\scriptscriptstyle\textup{{T}}}MG\right|\geq 1\right)$ for $M\in\mathbb{M}_{\mathrm{sym}}^{\ell}$ with ${\left\lVert M\right\rVert}_{\mathrm{op}}=1$. We use a simple form of a general technique due to [18, Lemma 1.2 ]. Let $\widetilde{G}=(I-vv^{\scriptscriptstyle\textup{{T}}})G+\widetilde{g}v$ where $v$ is the singular vector of $M$ corresponding to the singular value $1$, and $\widetilde{g}\sim\mathrm{Normal}(0,1)$ is independent of $G$. Since $G$ and $\widetilde{G}$ are both standard Gaussian random vectors, we have
@@
4.2 A concentration inequality for the sample covariance matrix of sub-exponential random vectors
[56, Theorem 3 ] also provides a concentration inequality for the sample covariance matrix of centered log-concave distributions. Following [2], [56] approximates the quadratic form induced by the sample covariance matrix by a term obtained through a certain direction-dependent truncation, and a term formed by the “peaky” residuals. The PAC-Bayesian argument is then used to obtain the tail bound for the “truncated” term, whereas the tail bound for peaky term is obtained using the decoupling-chaining method [49, Proposition 9.4.2] as cited in [56], combined with the tail bound for the Euclidean norm of log-concave random vectors due to [39, Thoerem 1.1 ] (see also [1, Theorem 2]). In fact, Paouris’s inequality is crucial in the proof of [56, Theorem 3 ] in the sense that it is tailored to log-concave measures instead of merely using the sub-exponential property. [^5] Log-concave measures are “regular” (see, e.g., [26, Definition 3.4], [29, the statement following Definition 2], [36]) which also implies that they are sub-exponential.
Our next result, Theorem˜4, provides a concentration inequality for the sample covariance matrix of random vectors with sub-exponential distribution. Similar to the model considered in Theorem˜1, the marginals of the sample covariance matrix do not have finite exponential moments, which rules out the analysis based solely on the PAC-Bayesian argument. While not adapted for log-concave distributions, without resorting to truncation techniques and relying only on Lemma˜1, we can partially recover the result of [56, Theorem 3 ] with a matching confidence level. Specifically, under the extra condition that the number of samples is at least the cube of the effective rank of the population covariance matrix, the tail bound we provide below in Theorem˜4 is equivalent to the tail bound provided by [56, Theorem 3 ].
@@ 4.2 A concentration inequality for the sample covariance matrix of sub-exponential random vectors
The proof is similar to the proof of Theorem˜3. Let $X=\sum_{i=1}^{n}\left(Z_{i}Z_{i}^{\scriptscriptstyle\textup{{T}}}-\Sigma\right)$ and $U=\kappa_{0}^{2}GG^{\scriptscriptstyle\textup{{T}}}$ for a parameter $\kappa_{0}>0$ and a standard Gaussian random vector $G\in\mathbb{R}^{d}$ that is independent of everything else. We can apply Lemma˜1 as in the proof of Theorem˜3; modification is only required to approximate the moments of $\left|{\langle U,X\rangle}\right|$. Using the standard symmetrization argument, for any $u\in\mathbb{R}^{d}$ we have
@@
where $\varepsilon_{1},\dotsc,\varepsilon_{n}$ are Rademacher random variables independent of each other and everything else. It then follows from [27, Corollary 2] that for $p\geq 2$ we have
@@
We can again show (23) still holds following the same argument used in the proof of Theorem˜3. Therefore, we have
@@
4.3 A tail bound for the operator norm of random matrix series
A classical model studied in random matrix theory is the random matrix series model
$X$ $=\sum_{i=1}^{n}\xi_{i}A_{i}\,.$ (27)
where the matrices $A_{1},\dotsc,A_{n}\in\mathbb{R}^{d_{1}\times d_{2}}$ are deterministic and the coefficients $\xi_{1},\dotsc,\xi_{n}$ are random variables. In particular, if $\xi_{i}$s are i.i.d. standard Gaussian (or Rademacher) random variables the non-commutative Khintchine inequality [32, 33] provides a bound on the Schatten-$q$ norms, including the operator norm, of $X$; in its prevalent formulation for $d\times d$ symmetric matrices $A_{i}$ (see, e.g., [50, Corollary 2.4]) it gurantees that
$\operatorname{\mathbb{E}}{\left\lVert X\right\rVert}_{\mathrm{op}}$ $\lesssim\sqrt{\log d}\,{\left\lVert\sum_{i=1}^{n}A_{i}^{2}\right\rVert}_{\mathrm{op}}^{1/2}\,.$
The suboptimal $O(\sqrt{\log(d)})$ multiplicative factor is eliminated recently in [9, 10] which, borrowing ideas from free probability, derived a sharp inequality by showing that the spectrum of $X$ is closely approximated by the spectrum of a certain corresponding infinite-dimensional deterministic operator $X_{\mathrm{free}}$. In particular, [10, Theorem 1.1 ] showed that
$\left|\operatorname{\mathbb{E}}\left({\left\lVert X\right\rVert}_{\mathrm{op}}\right)-{\left\lVert X_{\mathrm{free}}\right\rVert}\right|$ $\lesssim\widetilde{\upsilon}(X)(\log(d))^{3/4}\,$, (28)
where ${\left\lVert X_{\mathrm{free}}\right\rVert}\leq 2{\left\lVert\sum_{i\in[n]}A_{i}^{2}\right\rVert}_{\mathrm{op}}^{1/2}$ and $\widetilde{\upsilon}(X)={\left\lVert\sum_{i\in[n]}A_{i}^{2}\right\rVert}_{\mathrm{op}}^{1/4}\sup_{w\in\mathbb{R}^{n}\colon{}{\left\lVert w\right\rVert}_{2}\leq 1}{\left\lVert\sum_{i=1}^{n}w_{i}A_{i}\right\rVert}_{{\mathrm{F}}}^{1/2}$. [14] further showed the universality of these inequalities by considering $X$ to be a sum of independent random matrices and relating its spectrum to the spectrum of a Gaussian matrix whose mean and covariance operator match those of $X$.
The following theorem considers a more general model that accommodates the cases where the coefficients $(\xi_{i})_{i\in[n]}$ have heavier tail and perhaps are not even independent.
Theorem 5.
Let $\xi\in\mathbb{R}^{n}$ be a centered isotropic random vector such that for all $v\in\mathbb{R}^{n}$ we have
$\left(\operatorname{\mathbb{E}}\left(\left|{\langle\xi,v\rangle}\right|^{p}\right)\right)^{1/p}$ $\leq h(p)\left(\operatorname{\mathbb{E}}\left(|g|^{p}\right)\right)^{1/p}{\left\lVert v\right\rVert}_{2}\,$, (29)
where $g\sim\mathrm{Normal}(0,1)$ and the increasing function $h:[1,\infty]\mapsto\mathbb{R}_{>0}$ is a moment profile of $\xi$ relative to the standard Gaussian. Furthermore, given matrices $A_{1},\dotsc,A_{n}\in\mathbb{R}^{d_{1}\times d_{2}}$ let
$X$ $=\sum_{i=1}^{n}\xi_{i}A_{i}\,.$
For any $t\geq 0$, with probability at least $1-e^{-t}$ we have
${\left\lVert X\right\rVert}_{\mathrm{op}}$ $\lesssim\inf_{p\geq 2}e^{t/p}h(p)\left(\sigma_{*}p^{1/2}+\sigma+\upsilon+\sigma_{\diamond}p^{-1/2}\right)\,$, (30)
where
$\sigma_{*}$ $=\sup_{w\in\mathbb{R}^{n}\colon{}{\left\lVert w\right\rVert}_{2}\leq 1}{\left\lVert\sum_{i=1}^{n}w_{i}A_{i}\right\rVert}_{\mathrm{op}}\,$, $\upsilon$ $=\sup_{w\in\mathbb{R}^{n}\colon{}{\left\lVert w\right\rVert}_{2}\leq 1}{\left\lVert\sum_{i=1}^{n}w_{i}A_{i}\right\rVert}_{{\mathrm{F}}}\,$,
$\sigma$ $={\left\lVert\sum_{i=1}^{n}A_{i}^{\scriptscriptstyle\textup{{T}}}A_{i}\right\rVert}_{\mathrm{op}}^{1/2}+{\left\lVert\sum_{i=1}^{n}A_{i}A_{i}^{\scriptscriptstyle\textup{{T}}}\right\rVert}_{\mathrm{op}}^{1/2}$ $\sigma_{\diamond}$ $=\left(\sum_{i=1}^{n}{\left\lVert A_{i}\right\rVert}_{\mathrm{F}}^{2}\right)^{1/2}\,.$
In the special case of symmetric Gaussian series matrices (i.e., $\xi_{i}\overset{\mathrm{i.i.d.}}{\sim}\mathrm{Normal}(0,1)$) we have $h(p)=1$, and choosing $p=2t+2\sigma_{\diamond}/\sigma_{*}\geq 2$ in Theorem˜5 to obtain ${\left\lVert X\right\rVert}_{\mathrm{op}}\lesssim\sigma+\upsilon+\sigma_{\diamond}^{1/2}\sigma_{*}^{1/2}+\sqrt{2t}\sigma_{*}$ with probability at least $1-e^{-t}$. The bound
$\operatorname{\mathbb{E}}{\left\lVert X\right\rVert}_{\mathrm{op}}$ $\lesssim\sigma+\upsilon+\sigma_{\diamond}^{1/2}\sigma_{*}^{1/2}$ (31)
cab be extracted by integrating the tail bound (or directly applying the argument used to prove Lemma˜1). Using the PAC-Bayesian argument, [3, Theorem 1.2 ] obtained a bound for the expected injective norm [^6] For a tensor $T\in\mathbb{R}^{d_{1}\times\dotsb\times d_{k}}$ the injective norm of $T$ induced by the $q$-norm is ${\left\lVert T\right\rVert}_{\vee_{q}}=\sup\{\langle u_{1}\otimes\dotsb\otimes u_{k},T\rangle\colon{}\forall i\in[k]\ u_{i}\in\mathbb{R}^{d_{i}},\,{\left\lVert u_{i}\right\rVert}_{q}\leq 1\}$ of sub-Gaussian random tensor series (where $A_{i}$s are tensors and $\xi_{i}$s are sub-Gaussian). For matrices, the bound (30) and that of [3] are the same up to constant factors. The sub-optimal $\sqrt{\log d}$ factor of the non-commutative Khintchine inequality is eliminated in these bounds, but they are generally inferior to the sharp inequality (28). For example, for series with symmetric matrices we may have
$\sigma_{\diamond}\sigma_{*}=\left(\sum_{i=1}^{n}{\left\lVert A_{i}\right\rVert}_{{\mathrm{F}}}^{2}\right)^{1/2}\sup_{w\colon{\left\lVert w\right\rVert}_{2}\leq 1}{\left\lVert\sum_{i=1}^{n}w_{i}A_{i}\right\rVert}_{\mathrm{op}}$ $\gg{\left\lVert\sum_{i=1}^{n}A_{i}^{2}\right\rVert}_{\mathrm{op}}^{1/2}\sup_{w\colon{\left\lVert w\right\rVert}_{2}\leq 1}{\left\lVert\sum_{i=1}^{n}w_{i}A_{i}\right\rVert}_{{\mathrm{F}}}=\frac{\sigma\upsilon}{2}\,$,
where the left-hand side can be as large as $\sqrt{d}$ times the right-hand side (e.g., for $n=d$ and $A_{i}=e_{i}\otimes e_{i}$ for all $i\in[n]$).
The advantage of Theorem˜5 is that it addresses the situations where $\xi$ can have a general (relative) moment profile and does not impose the independence condition on the coordinates of $\xi$. For example, if $\xi$ is sub-exponential, even with independent coordinates, the standard PAC-Bayesian argument, as used in [3], cannot be employed since the analysis involves random variables that do not have finite exponential moments.
Proof of Theorem˜5.
For $i\in[2]$, let $G_{i}\in\mathbb{R}^{d_{i}}$ be independent standard Gaussian random vectors. Furthermore, for parameter $\kappa_{0}>0$, let $\operatorname{\mathbb{P}}_{0}$ to be the law of
$U$ $=\kappa_{0}G_{1}\otimes G_{2}\,.$
To apply Lemma˜1 we first find an upper bound for $\left(\operatorname{\mathbb{E}}_{0}\left(M_{X}(U,p)\right)\right)^{1/p}$. Let $X^{\natural}=\sum_{i\in[n]}\omega_{i}A_{i}$ where $\omega_{i}$s are i.i.d. standard Gaussian random variables. Observe that under the assumption (29) we have $M_{X}(u,p)\leq h^{p}(p)M_{X^{\natural}}(u,p)$ for every $u\in\mathbb{R}^{d_{1}\times d_{2}}$. Therefore, we have
$\left(\operatorname{\mathbb{E}}_{0}\left(M_{X}(U,p)\right)\right)^{1/p}$ $\leq h(p)\left(\operatorname{\mathbb{E}}_{0}\left(M_{X^{\natural}}(U,p)\right)\right)^{1/p}$
$=h(p)\left(\operatorname{\mathbb{E}}\left(\left|{\langle U,X^{\natural}\rangle}\right|^{p}\right)\right)^{1/p}\,.$
Let $A_{[n]}\in\mathbb{R}^{d_{1}\times d_{2}\times n}$ be the tensor formed by stacking $A_{i}$s along a third axis. With $G_{3}=\begin{bmatrix}\omega_{1},&\dotsb,&\omega_{n}\end{bmatrix}^{\scriptscriptstyle\textup{{T}}}$ and $p\geq 2$, applying the bound on the moments of Gaussian chaoses due to [28, Theorem 1 ] (see also [30]) to the Gaussian chaos ${\langle U,X^{\natural}\rangle}=\kappa_{0}{\langle G_{1}\otimes G_{2}\otimes G_{3},A_{[n]}\rangle}$ yields
$\left(\operatorname{\mathbb{E}}\left(\left|{\langle U,X^{\natural}\rangle}\right|^{p}\right)\right)^{1/p}$ $\lesssim\kappa_{0}\left(\sigma_{*}p^{3/2}+(\sigma+\upsilon)p+\sigma_{\diamond}p^{1/2}\right)\,.$
Therefore, we have
$\left(\operatorname{\mathbb{E}}_{0}\left(M_{X}(U,p)\right)\right)^{1/p}$ $\lesssim\kappa_{0}h(p)\left(\sigma_{*}p^{3/2}+(\sigma+\upsilon)p+\sigma_{\diamond}p^{1/2}\right)\,.$ (32)
Furthermore, for $x\in\mathbb{R}^{d_{1}\times d_{2}}$ with ${\left\lVert x\right\rVert}_{\mathrm{op}}=1$, let $(v_{i}\in\mathbb{R}^{d_{i}})_{i\in[2]}$ be the top singular vectors of $x$ for which ${\langle v_{1}\otimes v_{2},x\rangle}=1$. Also, for $i\in[2]$ let $\widetilde{G}_{i}=(I_{d_{i}}-v_{i}v_{i}^{\scriptscriptstyle\textup{{T}}})G_{i}+\widetilde{g}_{i}v_{i}$ where $I_{d_{i}}$ is the $d_{i}\times d_{i}$ identity matrix and $\widetilde{g}_{i}$ is a standard Gaussian random variable indepndent of everything else. Clearly, $\widetilde{G}_{i}\sim\mathrm{Normal}(0,I_{d_{i}})$ and $G_{i}-\widetilde{G}_{i}\stackrel{{\scriptstyle\textup{\tiny d}}}{{=}}\sqrt{2}g_{i}v_{i}$ where $\stackrel{{\scriptstyle\textup{\tiny d}}}{{=}}$ denotes equivalence in distribution and $g_{i}\sim\mathrm{Normal}(0,1)$. Let $Z_{1}=U$, $\widetilde{Z}_{1}=\kappa_{0}\widetilde{G}_{1}\otimes G_{2}$,
$Z_{2}$ $=\kappa_{0}(G_{1}-\widetilde{G}_{1})\otimes G_{2}\,$,
$\widetilde{Z}_{2}$ $=\kappa_{0}(G_{1}-\widetilde{G}_{1})\otimes\widetilde{G}_{2}\,$,
and
$Z_{3}$ $=\kappa_{0}(G_{1}-\widetilde{G}_{1})\otimes(G_{2}-\widetilde{G}_{2})$
$\stackrel{{\scriptstyle\textup{\tiny d}}}{{=}}2\kappa_{0}g_{1}g_{2}\,v_{1}\otimes v_{2}\,$,
and observe that $Z_{i+1}=Z_{i}-\widetilde{Z}_{i}$. For each $i\in[2]$ we have $Z_{i}\stackrel{{\scriptstyle\textup{\tiny d}}}{{=}}\widetilde{Z}_{i}$ and thereby
$2\operatorname{\mathbb{P}}\left(\left|{\langle Z_{i},x\rangle}\right|\geq 1\right)$ $=\operatorname{\mathbb{P}}\left(\left|{\langle Z_{i},x\rangle}\right|\geq 1\right)+\operatorname{\mathbb{P}}\left(\left|{\langle\widetilde{Z}_{i},x\rangle}\right|\geq 1\right)$
$\geq\operatorname{\mathbb{P}}\left(\left|{\langle Z_{i},x\rangle}\right|+\left|{\langle\widetilde{Z}_{i},x\rangle}\right|\geq 2\right)$
$\geq\operatorname{\mathbb{P}}\left(\left|{\langle Z_{i+1},x\rangle}\right|\geq 2\right)\,.$
Therefore, we obtain
$\operatorname{\mathbb{P}}\left(\left|{\langle Z_{1},x\rangle}\right|\geq 1\right)$ $\geq 4^{-1}\operatorname{\mathbb{P}}\left(\left|{\langle Z_{3},x\rangle}\right|\geq 4\right)\,.$
Recalling the definitions of $Z_{1}$ and $Z_{3}$, and the fact that ${\langle v_{1}\otimes v_{2},x\rangle}={\left\lVert x\right\rVert}_{\mathrm{op}}=1$ by choice, it follows that
$\operatorname{\mathbb{P}}_{0}\left(\left|{\langle U,x\rangle}\right|\geq 1\right)$ $\geq 4^{-1}\operatorname{\mathbb{P}}\left(\kappa_{0}\left|g_{1}g_{2}\right|\geq 2\right)$
$\geq 4^{-1}\operatorname{\mathbb{P}}\left(|g_{i}|\geq\sqrt{2}\kappa_{0}^{-1/2}\ \text{for all}\ i\in[2]\right)$
$=\prod_{i\in[2]}\Phi(-\sqrt{2}\kappa_{0}^{-1/2})$
$\geq 4^{-2}e^{-8/(3\kappa_{0})}\,$,
where again we used (12) in the fourth line. Consequently $\nu(\operatorname{\mathbb{P}}_{0})\leq 16e^{8/(3\kappa_{0})}$. In view of (32), it follows from Lemma˜1 that with probability at least $1-e^{-t}$ we have
${\left\lVert X\right\rVert}_{\vee}$ $\lesssim\inf_{p\geq 2,\,\kappa_{0}>0}e^{t/p}16^{1/p}e^{8/(3p\kappa_{0})}\kappa_{0}h(p)\left(\sigma_{*}p^{3/2}+(\sigma+\upsilon)p+\sigma_{\diamond}p^{1/2}\right)\,.$
Simple calculus shows that the choice of $\kappa_{0}=8/(3p)$ is optimal. Evaluating the objective at this value yields (30). ∎